OpenStudy (anonymous):

The x-intercept of the line is -5 and the y-intercept of the line -2. What is the equation of the line? y = -5/2x - 5 y = 2/5x + 2 y = 5/2x - 2 y = -2/5x - 2

OpenStudy (like1234):

is there a picture

OpenStudy (anonymous):

No

OpenStudy (texaschic101):

in y = mx + b form, the y intercept is in the b position...so if the y intercept is -2, then we know, so far, that the equation will be set up like : y = mx - 2. So that eliminated the first 2 answer choices. Now...to find the x intercept, sub in 0 for y... y = 5/2x - 2 0 = 5/2x - 2 -5/2x = -2 x = -2/(-5/2) x = -2 * -2/5 x = 4/5.....not this one y = -2/5x - 2 0 = -2/5x - 2 2/5x = -2 x = -2/(2/5) x = -2 * 5/2 x = - 5 ------so this one has a y intercept of -2 and an x intercept of -5 conclusion : answer is y = -2/5x - 2 any questions ?

OpenStudy (anonymous):

What is the slope-intercept form of the equation y - 7 = -5/2 (x + 4)? y - 2 = -5/2 (x + 2) y + 7 = -x + 5/2 y = -4/7 + 2 y = -5/2 x - 3

OpenStudy (anonymous):

@texaschic101